# The Interplay of Heterogeneity and Product Detachment in Templated Polymer Copying

### v1

@@ Abstract
Templated copolymerization, in which information stored in the sequence of a heteropolymer template is copied into another polymer product, is the mechanism behind all known methods of genetic information transfer. A key aspect of templated copolymerization is the eventual detachment of the product from the template. A second key feature of natural biochemical systems is that the template-binding free energies of both correctly-matched and incorrect monomers are heterogeneous. Previous work has considered the thermodynamic consequences of detachment, and the consequences of heterogeniety for polymerisation speed and accuracy, but the interplay of both separation and heterogeniety remains unexplored. In this work, we investigate a minimal model of templated copying that simultaneously incorporates both detachment from behind the leading edge of the growing copy and heterogeneous interactions. We first extend existing coarse-graining methods for models of polymerisation to allow for heterogeneous interactions. We then show that heterogeneous copying systems with explicit detachment do not exhibit the subdiffusive behaviour observed in the absence of detachment when near equilibrium. Next, we show that heterogeneity in correct monomer interactions tends to result in slower, less accurate copying, while heterogeneity in incorrect monomer interactions tends to result in faster, more accurate copying, due to an increased roughness in the free energy landscape of either correct or incorrect monomer pairs. Finally, we show that heterogeneity can improve on known thermodynamic efficiencies of homogeneous copying, but these increased thermodynamic efficiencies do not always translate to increased efficiencies of information transfer.
@@ I Introduction
More detailed models have also been made that focus on the polymerisation mechanism, typically discrete state Markov models. The simplest ones involve one step per added monomer [28, 29, 30, 31, 32, 33, 34]; more complex ones have multiple steps [35] and sometimes added cycles [36, 37, 38, 39, 40, 41, 42, 43]. Some models explicitly consider microscopic reversibility [28, 44, 29, 30, 31, 32, 33, 34, 43], while others incorporate irreversible steps [36, 37, 38, 40, 41, 42]. Another distinction is whether they assume a symmetry between all correct pairings and, separately, all incorrect pairings (homogeneous) [36, 37, 28, 44, 29, 39, 40, 38, 31, 41, 42, 33, 35, 43] or consider explicit differences between them (heterogeneous) [32, 34]. Finally, some approaches account for continuous separation of the product from the template [29, 33, 35, 43], or just model assembly on the template [36, 37, 28, 44, 39, 40, 38, 30, 31, 41, 42, 34].
@@
To date, heterogeneity has not been considered in the context of templated copying systems that incorporate explicit product detachment. The effects of heterogeneity in models without detachment have been shown to be non-trivial because each site on the template does not contribute independently to the overall growth dynamics, resulting in correlations within the product and the emergence of a rugged landscape with significant effects on copying velocities [44, 30, 32]. It is unclear if and how phenomena observed in heterogeneous, non-separating systems translate to systems that incorporate detachment. We fill this gap by incorporating heterogeneity into a known model of polymer copying that incorporates separation from behind the growing tip of the product, analogous to transcription and translation [33]. The addition of each monomer in our model consists of multiple steps, following [35]. We first introduce a method of coarse-graining the model to a single-step description, allowing existing methods ([32] and [43]) of analysis to be used. Using these methods, we demonstrate qualitative differences between heterogeneous copying with and without separation. We then identify regimes where heterogeneity facilitates or hinders fast, accurate and efficient copying.
@@
We organize our work as follows. In Sections II.1 and II.2, we introduce our coarse-grained and fine-grained Markov copying models with separation. We briefly describe Gaspard’s and Qureshi’s analysis methods in Section II.3 and our extended iteration for state visits in Section II.4. In Section II.5, we describe the parameter sets we initially consider. We begin our results in Section III.1 by comparing velocities of heterogeneous copying with and without separation. In Section III.2, we investigate the effect of heterogeneity on error and completion time for our limiting regimes. Finally, in Section III.3, we show that in the low driving regime, the thermodynamic efficiency of copying can be improved by heterogeneity.
@@ II.1 General Coarse-Grained Model of Templated Polymer Copying with Heterogeneity and Product Separation
Figure 1: Polymer copying with continuous separation of the product from the template. (a) Illustration of the coarse-grained model for a two-monomer system. A coarse-grained state is represented by the full sequence of the copy and the template. At each step, one of either type of monomer may be added onto the growing copy sequence, or the existing tip monomer in the copy may be removed. Each coarse-grained step (purple dots) may be further subdivided into fine-grained steps. (b) Illustration of the fine-grained steps that occur between coarse-grained states. Specifically, the fine-grained steps corresponding to the leftmost purple dot in subfigure (a) are depicted. First, a new monomer binds to the template with both a generic (dashed line) and sequence-specific (solid line) bond. Then, the newly bound monomer is covalently incorporated into the growing polymer, breaking the generic bond in the process. Finally, the previously added monomer unbinds from the template. (c) Labelling of fine-grained states for a given coarse-grained state, with indices $f$ denoted by bold bracketed numbers. Transitions to other states of index $f=0$ (denoted here by the dotted arrows) change the coarse-grained state $(\mathbf{x},\mathbf{y})$. Note that these fine-grained state indices $f$ are not valid when returning from a different coarse-grained state (via a dashed arrow) as $f$ is dependent on the last completed state visited. For simplicity, general and sequence-specific interactions are not depicted separately in subfigures (a) and (c)
@@ II.2 Fine-Grained Reaction Steps
The monomer addition and removal steps in the previous section are not elementary chemical reactions, and in a general model will be described through a series of smaller, fine-grained steps. Properties of the fine-grained system can be preserved in the coarse-grained system if an appropriate coarse-graining procedure is followed (we cover this procedure for our case in Sections II.3 and II.4; refer to Qureshi [43] for more detail). Fine-grained steps for our system follow from earlier work [35] on a copying system that separates as it is being copied; in brief, monomer binding, polymerization and unbinding of the copy from the template are each treated as separate steps. Note that although we use the term fine-grained, these states should still be understood as biochemical macrostates [45].
@@
We use the notation $(\mathbf{x},\mathbf{y},f)$ to specify the fine-grained states, consisting of the coarse-grained state as well as an index $f\in\{0,..,6\}$ that specifies the fine-grained state iven the coarse-grained state (Figure 1(c)). We treat the fine-grained process as a series of sub-processes in which the system transitions between “completed states” with $f=0$, via “transitory states” with $f\neq 0$, as in [43]. The coarse-grained label $(\mathbf{x},\mathbf{y})$ is given by the completed state that was visited most recently, with $f$ then defined relative to that state. Hence, biochemical macrostates with $f\neq 0$ have two different fine-grained state descriptors $\mathbf{y}$ and $f$ depending on whether they were approached “from the back” or “from the front”. For example, starting from $(\mathbf{x},\&m_{l},0)$, binding and then polymerization of a type $1$ monomer would yield $(\mathbf{x},\&m_{l},3)$. On the other hand, the same state would be labelled $(\mathbf{x},\&m_{l}1,1)$ if $(\mathbf{x},\&m_{l}1,0)$ was the last completed state visited.
@@ II.3 Solving for Distributions of Complete Polymers
There are many alternative ways that $S^{\mathbf{x}}_{F}(t)$ can be coarse-grained with a view to avoiding the need of handling non-Markovian processes. Qureshi’s coarse-graining procedure [43] can be applied if the full state space of the system can be partitioned into transitory and completed states such that every transitory state is only encountered during a transition between two unique completed states. In such systems, between any two completed states is a ‘petal’ of transitory states. Each petal between completed states can then be analyzed independently.
For our model, transitory states $(\mathbf{x},\&m_{l}m_{l+1},f)$ for $f\in\{1,2\}$ (Figure 1(c)) are accessible only during transitions from completed state $(\mathbf{x},\&m_{l}m_{l+1},0)$ to $(\mathbf{x},\&m_{l},0)$. All other transitory states are accessible only during transitions from $(\mathbf{x},\&m_{l}m_{l+1},0)$ to $(\mathbf{x},\&m_{l}m_{l+1}m_{l+2},0)$ for some $m_{l+2}$. Hence, the entire set of states and transitions in Figure 1(c), which connect two completed states together, is a ‘petal’ that can be analyzed independently of others. Qureshi’s procedure is based on a graphical interpretation of discrete-state markov processes, where states in a petal are nodes in a directed weighted graph and transitions between states are directed edges in the graph with weights equal to transition rates. Qureshi then considers rooted spanning trees of this graph for a single petal - subgraphs without cycles where the out-degree of every node is equal to 1 except for the root node, which has out-degree zero (we say the spanning tree is “rooted at” the root node). Given a starting state $(\mathbf{x},\mathbf{y,0})$ and target state $(\mathbf{x},\mathbf{y}^{*},0)$, the key quantity $\Lambda^{+}_{\mathbf{x}}(\mathbf{y,y^{*}})$ can be calculated by multiplying the rates of all transitions in each spanning tree rooted at $(\mathbf{x},\mathbf{y}^{*},0)$ and then summing over these spanning trees. $\Lambda^{-}_{\mathbf{x}}(\mathbf{y,y^{*}})$ can be obtained by doing the same over spanning trees rooted at $(\mathbf{x},\mathbf{y},0)$. Summing over spanning trees rooted at $(\mathbf{x},\mathbf{y},0)$ for a modified process where all transitions to $(\mathbf{x},\mathbf{y}^{*},0)$ are redirected back to $(\mathbf{x},\mathbf{y},0)$ yields a second key quantity $A_{\mathbf{x}}(\mathbf{y,y^{*}})$. Qureshi’s procedure assigns the coarse-grained backward and forward propensities from $\mathbf{y}$ to $\mathbf{y^{*}}$ as follows:
@@
This coarse-graining procedure, yielding both the Markov process $S^{\mathbf{x}}_{C}(t)=(Y_{C}(t)|X=\mathbf{x})$ and the embedded sequence of states $S^{\mathbf{x}}_{C}[k]$, has two advantages. First, the resultant addition and removal rates obey the constraint $\Phi^{+}(m_{l+1},\mathbf{x},\mathbf{y})=\Phi^{+}_{l}(\mathbf{x},m_{l}m_{l+1})$ and $\Phi^{-}(\mathbf{x},\mathbf{y})=\Phi^{-}_{l}(\mathbf{x},m_{l-1}m_{l})$, facilitating the use of Gaspard’s iterated function system, which we will describe shortly. Second, it preserves the distribution of the sequence of coarse-grained states visited: $S^{\mathbf{x}}_{C}[k]=\tilde{S}^{\mathbf{x}}_{F}[k]$. Hence, if we are interested in the distribution of complete polymers (defined as the distribution of polymers obtained after complete detachment from the template) and derivative quantities like error rate, then we can work with the Markovian coarse-grained processes $S^{\mathbf{x}}_{C}(t)$ and $S^{\mathbf{x}}_{C}[k]$, while completely ignoring the fine-grained steps other than when calculating $\Lambda_{\mathbf{x}}^{\pm}(\mathbf{y,y^{*}})$ and $A_{\mathbf{x}}(\mathbf{y,y^{*}})$. However, in general $S^{\bf x}_{C}(t)\neq\tilde{S}^{\bf x}_{F}(t)$, which means that polymer completion times and velocities are not preserved by this mapping. In the presence of cycles in the fine-grained steps, free-energy consumption will also not be preserved (this fact is less important for us as our fine-grained system has no cycles). Qureshi describes a method for recovering these properties in the homogeneous case [43], but we need to derive an additional result (Section II.4) in order to adapt his method to the heterogeneous case.
@@
Figure 2: Markov chain first-passage problem for the calculation of complete polymer distributions with a length $l$ product polymer as an initial state. The $p_{\mathbf{x},l}(m_{l+1}|m_{l-1}m_{l})$ variables are local transition probabilities that can be obtained by considering coarse-grained propensities $\Phi^{+}_{l}$ or $\Phi^{-}_{l}$. The $Q$ variables are local probabilities of absorption (into some arbitrary complete polymer) occurring before monomer removal. The absorption probabilities $R_{\mathbf{x},l}(m_{l+1}|m_{l-1}m_{l})$ follow from values of $Q$ for the next position index $Q_{\mathbf{x},l+1}(m_{l}m_{l+1})$. The $Q$ values (since these are the probabilities of absorption before removal) for the current iteration step can then be calculated by summing over all the non-backward absorption probabilities, in this case $Q_{\mathbf{x},l}(21)=R_{\mathbf{x},l}(1|21)+R_{\mathbf{x},l}(2|21)$. With this current $Q$ value, the next step of the backward iteration may proceed.
@@
Having argued that quantities of the fine-grained model can be obtained from solutions of a Markovian coarse-grained model, we now proceed to discuss the method we use to actually solve for quantities of the coarse-grained model. Gaspard showed that the distribution of consecutive copy monomers, conditioned on the entire template, is Markovian if transition rates are only dependent on tip states [30, 32]. Moreover, the conditional probability distributions $P_{\mathbf{x},l}(m_{l}|m_{l-1})$ for complete polymers given a fixed template $\mathbf{x}$ and position indices $l$ can be obtained through the solution of an iterated function system. We discuss here a brief intuition for Gaspard’s iterated function system in terms of absorbing probabilities of Markov chains, and perform a full derivation for our system in Appendix A (refer to [32] for a full treatment).

### v2

@@ Abstract
Templated copolymerization, in which information stored in the sequence of a heteropolymer template is copied into another polymer product, is the mechanism behind all known methods of genetic information transfer. A key aspect of templated copolymerization is the eventual detachment of the product from the template. A second key feature of natural biochemical systems is that the template-binding free energies of both correctly-matched and incorrect monomers are heterogeneous. Previous work has considered the thermodynamic consequences of detachment, and the consequences of heterogeneity for polymerisation speed and accuracy, but the interplay of both separation and heterogeneity remains unexplored. In this work, we investigate a minimal model of templated copying that simultaneously incorporates both detachment from behind the leading edge of the growing copy and heterogeneous interactions. We first extend existing coarse-graining methods for models of polymerisation to allow for heterogeneous interactions. We then show that heterogeneous copying systems with explicit detachment do not exhibit the subdiffusive behaviour observed in the absence of detachment when near equilibrium. Next, we show that heterogeneity in correct monomer interactions tends to result in slower, less accurate copying, while heterogeneity in incorrect monomer interactions tends to result in faster, more accurate copying, due to an increased roughness in the free energy landscape of either correct or incorrect monomer pairs. Finally, we show that heterogeneity can improve on known thermodynamic efficiencies of homogeneous copying, but these increased thermodynamic efficiencies do not always translate to increased efficiencies of information transfer.
@@ I Introduction
More detailed models have also been made that focus on the polymerisation mechanism, typically discrete state Markov models. The simplest ones involve one step per added monomer [28, 29, 30, 31, 32, 33, 34]; more complex ones have multiple steps [35] and sometimes added cycles [36, 37, 38, 39, 40, 41, 42, 43]. Some models explicitly consider microscopic reversibility [28, 44, 29, 30, 31, 32, 33, 34, 43], while others incorporate irreversible steps [36, 37, 38, 40, 41, 42]. Another distinction is whether they assume a symmetry between all correct pairings and, separately, all incorrect pairings (homogeneous) [36, 37, 28, 44, 29, 39, 40, 38, 31, 41, 42, 33, 35, 43] or consider explicit differences between them (heterogeneous) [32, 34]. Finally, some approaches are not preoccupied with copy-template separation [36, 37, 28, 44, 39, 40, 38, 30, 31, 42, 34]. On the other hand, other works consider models in which the copy explicitly separates from the template [41, 45], including some that attempt to understand the thermodynamic consequences of separation [29, 33, 35, 43].
@@
Although there are multiple theoretical frameworks to treat heterogeneity in the context of separating copolymerization systems [45, 32, 46, 47] (in particular, [45] explicitly applies their framework to RNA polymerase), existing work has not considered both separation and heterogeneity in models with thermodynamic self-consistency. Thermodynamically self-consistent models have been constructed for homogeneous systems [29, 33, 35, 43], but it remains unclear if and how the effects of heterogeneity interplay with the thermodynamic consequences of separation. We fill this gap by incorporating heterogeneity into a known thermodynamically consistent model of polymer copying that incorporates separation from behind the growing tip of the product, analogous to transcription and translation [33]. The addition of each monomer in our model consists of multiple steps, following [35]. We first introduce a method of coarse-graining the model to a single-step description, allowing existing methods ([32] and [43]) of analysis to be used. Using these methods, we demonstrate qualitative differences between heterogeneous copying with and without separation. We then identify regimes where heterogeneity facilitates or hinders fast, accurate and efficient copying.
@@
We organize our work as follows. In Sections II.1 and II.2, we introduce our coarse-grained and fine-grained Markov copying models with separation. We briefly describe Gaspard’s [32] and Qureshi et al’s [43] analysis methods in Section II.3 and our extended iteration for state visits in Section II.4. In Section II.5, we describe the parameter sets we initially consider. We begin our results in Section III.1 by comparing velocities of heterogeneous copying with and without separation. In Section III.2, we investigate the effect of heterogeneity on error and completion time for our limiting regimes. Finally, in Section III.3, we show that in the low driving regime, the thermodynamic efficiency of copying can be improved by heterogeneity.
@@ II.1 General Coarse-Grained Model of Templated Polymer Copying with Heterogeneity and Product Separation
Figure 1: Polymer copying with continuous separation of the product from the template. (a) Illustration of the coarse-grained model for a two-monomer system. Each monomer type is assigned a unique number (1 or 2 in this case) and color (light blue for 1 and red for 2). Square monomers correspond to the template polymer, while round monomers correspond to the copy polymer. A coarse-grained state is represented by the full sequence of the copy and the template polymers. Purple dotted arrows represent coarse-grained state transitions corresponding to either monomer addition (forward arrow) or monomer removal (backward arrow). Each coarse-grained transition may be further subdivided into fine-grained steps. (b) Illustration of the fine-grained steps that occur between coarse-grained states. Specifically, the fine-grained steps corresponding to the leftmost purple dot in subfigure (a) are depicted (only monomers in the final 3 positions of the copy and template are depicted). First, a new monomer binds to the template with both a generic (dashed line) and sequence-specific (solid line) bond. Then, the newly bound monomer is covalently incorporated into the growing polymer, breaking the generic bond in the process. Finally, the previously added monomer unbinds from the template. (c) Labelling of fine-grained states for a given coarse-grained state, with indices $f$ denoted by bold bracketed numbers. Arrows here correspond to fine-grained state transitions. Transitions to other states of index $f=0$ (denoted here by the dotted arrows) change the coarse-grained state $(\mathbf{x},\mathbf{y})$. Note that these fine-grained state indices $f$ are not valid when returning from a different coarse-grained state (via a dashed arrow) as $f$ is dependent on the last completed state visited. For simplicity, general and sequence-specific interactions are not depicted separately in subfigures (a) and (c)
@@ II.2 Fine-Grained Reaction Steps
The monomer addition and removal steps in the previous section are not elementary chemical reactions, and in a general model will be described through a series of smaller, fine-grained steps. Properties of the fine-grained system can be preserved in the coarse-grained system if an appropriate coarse-graining procedure is followed (we cover this procedure for our case in Sections II.3 and II.4; refer to Qureshi et al. [43] for more detail). Fine-grained steps for our system follow from earlier work [35] on a copying system that separates as it is being copied; in brief, monomer binding, polymerization and unbinding of the copy from the template are each treated as separate steps. Note that although we use the term fine-grained, these states should still be understood as biochemical macrostates [48].
@@
We use the notation $(\mathbf{x},\mathbf{y},f)$ to specify the fine-grained states, consisting of the coarse-grained state as well as an index $f\in\{0,..,6\}$ that specifies the fine-grained state given the coarse-grained state (Figure 1(c)). We treat the fine-grained process as a series of sub-processes in which the system transitions between “completed states” with $f=0$, via “transitory states” with $f\neq 0$, as in [43]. The coarse-grained label $(\mathbf{x},\mathbf{y})$ is given by the completed state that was visited most recently, with $f$ then defined relative to that state. Hence, biochemical macrostates with $f\neq 0$ have two different fine-grained state descriptors $\mathbf{y}$ and $f$ depending on whether they were approached “from the back” or “from the front”. For example, starting from $(\mathbf{x},\&m_{l},0)$, binding and then polymerization of a type $1$ monomer would yield $(\mathbf{x},\&m_{l},3)$. On the other hand, the same state would be labelled $(\mathbf{x},\&m_{l}1,1)$ if $(\mathbf{x},\&m_{l}1,0)$ was the last completed state visited.
@@ II.3 Solving for Distributions of Complete Polymers
There are many alternative ways that $S^{\mathbf{x}}_{F}(t)$ can be coarse-grained with a view to avoiding the need of handling non-Markovian processes. Qureshi et al.’s coarse-graining procedure [43] can be applied if the full state space of the system can be partitioned into transitory and completed states such that every transitory state is only encountered during a transition between two unique completed states. In such systems, between any two completed states is a ‘petal’ of transitory states. Each petal between completed states can then be analyzed independently.
For our model, transitory states $(\mathbf{x},\&m_{l}m_{l+1},f)$ for $f\in\{1,2\}$ (Figure 1(c)) are accessible only during transitions from completed state $(\mathbf{x},\&m_{l}m_{l+1},0)$ to $(\mathbf{x},\&m_{l},0)$. All other transitory states are accessible only during transitions from $(\mathbf{x},\&m_{l}m_{l+1},0)$ to $(\mathbf{x},\&m_{l}m_{l+1}m_{l+2},0)$ for some $m_{l+2}$. Hence, the entire set of states and transitions in Figure 1(c), which connect two completed states together, is a ‘petal’ that can be analyzed independently of others. Qureshi et al.’s procedure [43] is based on a graphical interpretation of discrete-state markov processes, where states in a petal are nodes in a directed weighted graph and transitions between states are directed edges in the graph with weights equal to transition rates. Qureshi et al. then consider rooted spanning trees of this graph for a single petal - subgraphs without cycles where the out-degree of every node is equal to 1 except for the root node, which has out-degree zero (we say the spanning tree is “rooted at” the root node). Given a starting state $(\mathbf{x},\mathbf{y,0})$ and target state $(\mathbf{x},\mathbf{y}^{*},0)$, the key quantity $\Lambda^{+}_{\mathbf{x}}(\mathbf{y,y^{*}})$ can be calculated by multiplying the rates of all transitions in each spanning tree rooted at $(\mathbf{x},\mathbf{y}^{*},0)$ and then summing over these spanning trees. $\Lambda^{-}_{\mathbf{x}}(\mathbf{y,y^{*}})$ can be obtained by doing the same over spanning trees rooted at $(\mathbf{x},\mathbf{y},0)$. Summing over spanning trees rooted at $(\mathbf{x},\mathbf{y},0)$ for a modified process where all transitions to $(\mathbf{x},\mathbf{y}^{*},0)$ are redirected back to $(\mathbf{x},\mathbf{y},0)$ yields a second key quantity $A_{\mathbf{x}}(\mathbf{y,y^{*}})$. Qureshi et al.’s procedure [43] assigns the coarse-grained backward and forward propensities from $\mathbf{y}$ to $\mathbf{y^{*}}$ as follows:
@@
This coarse-graining procedure, yielding both the Markov process $S^{\mathbf{x}}_{C}(t)=(Y_{C}(t)|X=\mathbf{x})$ and the embedded sequence of states $S^{\mathbf{x}}_{C}[k]$, has two advantages. First, the resultant addition and removal rates obey the constraint $\Phi^{+}(m_{l+1},\mathbf{x},\mathbf{y})=\Phi^{+}_{l}(\mathbf{x},m_{l}m_{l+1})$ and $\Phi^{-}(\mathbf{x},\mathbf{y})=\Phi^{-}_{l}(\mathbf{x},m_{l-1}m_{l})$, facilitating the use of Gaspard’s iterated function system [32], which we will describe shortly. Second, it preserves the distribution of the sequence of coarse-grained states visited: $S^{\mathbf{x}}_{C}[k]=\tilde{S}^{\mathbf{x}}_{F}[k]$. Hence, if we are interested in the distribution of complete polymers (defined as the distribution of polymers obtained after complete detachment from the template) and derivative quantities like error rate, then we can work with the Markovian coarse-grained processes $S^{\mathbf{x}}_{C}(t)$ and $S^{\mathbf{x}}_{C}[k]$, while completely ignoring the fine-grained steps other than when calculating $\Lambda_{\mathbf{x}}^{\pm}(\mathbf{y,y^{*}})$ and $A_{\mathbf{x}}(\mathbf{y,y^{*}})$. However, in general $S^{\bf x}_{C}(t)\neq\tilde{S}^{\bf x}_{F}(t)$, which means that polymer completion times and velocities are not preserved by this mapping. In the presence of cycles in the fine-grained steps, free-energy consumption will also not be preserved (this fact is less important for us as our fine-grained system has no cycles). Qureshi et al. describes a method for recovering these properties in the homogeneous case [43], but we need to derive an additional result (Section II.4) in order to adapt his method to the heterogeneous case.
@@
Figure 2: Markov chain absorption probability problem for the calculation of complete polymer distributions with a length $l$ product polymer as an initial state. Single arrows (without corresponding reverse counterparts) represent transitions to absorbing states (monomer removal to a copy of length $l-1$ and copy completion to some complete polymer for some given moomer at position $l+1$), and the probability of absorption to each absorbing state is considered. The length of the copy for each state is given at the top of the figure (note the jump from l+1 to L, indicating absorption into a complete and fully detached polymer). The $p_{\mathbf{x},l}(m_{l+1}|m_{l-1}m_{l})$ variables are local transition probabilities that can be obtained by considering coarse-grained propensities $\Phi^{+}_{l}$ or $\Phi^{-}_{l}$. The $Q$ variables are local probabilities of absorption (into some arbitrary complete polymer) occurring before monomer removal. The absorption probabilities $R_{\mathbf{x},l}(m_{l+1}|m_{l-1}m_{l})$ follow from values of $Q$ for the next position index $Q_{\mathbf{x},l+1}(m_{l}m_{l+1})$. The $Q$ values (since these are the probabilities of absorption before removal) for the current iteration step can then be calculated by summing over all the non-backward absorption probabilities, in this case $Q_{\mathbf{x},l}(21)=R_{\mathbf{x},l}(1|21)+R_{\mathbf{x},l}(2|21)$. With this current $Q$ value, the next step of the backward iteration may proceed.
@@
Having argued that quantities of the fine-grained model can be obtained from solutions of a Markovian coarse-grained model, we now proceed to discuss the method we use to actually solve for quantities of the coarse-grained model. Gaspard and Andrieux showed that the distribution of consecutive copy monomers, conditioned on the entire template, is Markovian if transition rates are only dependent on tip states [30, 32]. Moreover, the conditional probability distributions $P_{\mathbf{x},l}(m_{l}|m_{l-1})$ for complete polymers given a fixed template $\mathbf{x}$ and position indices $l$ can be obtained through the solution of an iterated function system. We discuss here a brief intuition for Gaspard’s iterated function system [32] in terms of absorbing probabilities of Markov chains, and perform a full derivation for our system in Appendix A (refer to [32] for a full treatment).